100ml of water sample on titration with N/50 H2SO4 required 12.4ml of the acid


for neutralization to phenolphthalein end point and 15.2ml of the acid to methyl


end endpoint. Determine the type and extent of alkalinity present in the water


sample.

1
Expert's answer
2022-01-21T03:49:02-0500

"Alkalinity=\frac{A*N*50000}{ml\>of\>sample}"


Phenolphthalein Alkalinity"=\frac{12.4×0.02×50000}{109}=124ppm"


Methylorange alkalinity

"=\frac{25.6×0.02×50000}{100}=156ppm"


The alkalinity is due to OH- and CO32- ions


Half of CO32- ions "=156-124=28ppm"


OH- Alkalinity "=124-28=96ppm"

"CO_3^{2-}" Alkalinity "=28\u00d72"

"=56ppm"